# Towards Scalable and Stable Parallelization of Nonlinear RNNs

### v1

@@ Abstract
Conventional nonlinear RNNs are not naturally parallelizable across the sequence length, whereas transformers and linear RNNs are. Lim et al. (2024) therefore tackle parallelized evaluation of nonlinear RNNs by posing it as a fixed point problem, solved with Newton’s method. By deriving and applying a parallelized form of Newton’s method, they achieve huge speedups over sequential evaluation. However, their approach inherits cubic computational complexity and numerical instability. We tackle these weaknesses. To reduce the computational complexity, we apply quasi-Newton approximations and show they converge comparably to full-Newton, use less memory, and are faster. To stabilize Newton’s method, we leverage a connection between Newton’s method damped with trust regions and Kalman smoothing. This connection allows us to stabilize Newtons method, per the trust region, while using efficient parallelized Kalman algorithms to retain performance. We compare these methods empirically, and highlight the use cases where each algorithm excels.
1 Statistics Department, Stanford University, Stanford, CA, USA
2 Wu Tsai Neurosciences Institute, Stanford University, Stanford, CA, USA
3 GE Healthcare, Seattle, WA, USA
4 Institute for Computational and Mathematical Engineering, Stanford University, Stanford, CA, USA
5 Liquid AI, Palo Alto, CA, USA
Correspondence should be addressed to X.G. (xavier18@stanford.edu) or S.W.L. (scott.linderman@stanford.edu)
@@ 1 Introduction
Parallel computation has helped fuel the rise of deep learning (Hooker, 2021). Architectures such as transformers (Vaswani et al., 2017) and linear RNNs (Katharopoulos et al., 2020; Smith et al., 2023; Orvieto et al., 2023; Gu and Dao, 2023) are specifically designed to allow parallelization over the length of the input sequence. However, most conventional nonlinear RNNs (e.g. Elman RNNs, GRUs (Cho et al., 2014), LSTMS (Hochreiter and Schmidhuber, 1997) etc.) are not readily parallelizable over the sequence length due to their inherently sequential architecture, and thus they do not benefit as greatly from parallel hardware. Nonetheless, these nonlinear RNN architectures are still used widely across the scientific community. Furthermore, recent work has suggested that linear RNNs (and transformers) are fundamentally limited in their expressivity compared to nonlinear RNNs (Merrill et al., 2024). Finally, nonlinear RNNs continue to be of significant interest in computational and theoretical neuroscience as models of neural systems (Sussillo and Barak, 2013; Smith et al., 2021). Therefore, scalable and stable parallelization methods for nonlinear RNNs offer significant benefits across many fields.
@@
Towards this goal, Lim et al. (2024) proposed DEER, a method for evaluating a nonlinear RNN in parallel by considering the hidden states as the solution of a fixed point equation. The particular fixed point equation enforces that the solution satisfies the nonlinear dynamics of the RNN. Newton’s method is used to solve the resulting fixed point equations with quadratic convergence rates (Nocedal and Wright, 2006, Chapter 11). Lim et al. (2024) also show that the inversion of the structured Jacobian matrix required by Newton’s method can be cast as an associative parallel scan (Blelloch, 1990). DEER therefore reduces the evaluation runtime over sequential evaluation by as much as factor of twenty.
@@
However, DEER also inherits the weaknesses of Newton’s method and parallel scans. The first weakness is scalability. Let $D$ denote the state dimension and $T$ denote sequence length. From using a parallel scan to evaluate updates from Newton’s method, DEER inherits $\mathcal{O}(TD^{2})$ memory complexity and $\mathcal{O}(TD^{3})$ computational work (Blelloch, 1990). These costs can be prohibitive in practical deep learning settings. The second limitation of DEER is numerical stability, inherited from Newton’s method. In general, undamped Newton’s method does not provide global convergence guarantees, and in practice often diverges (Nocedal and Wright, 2006).
We seek to ameliorate these weaknesses. In Section 3.2, we prove that DEER (under relatively weak assumptions) theoretically converges globally for RNNs; and show that in practice a simple heuristic recovers this behavior at the expense of increased runtime. We then use intuitions from this result in Section 4.1 to improve the scalability of DEER by applying quasi-Newton approximations. We refer to this quasi-Newton algorithm as quasi-DEER. We also improve the numerical stability of DEER by adding a trust region (Conn et al., 2000). By leveraging a connection between Newton’s method with a trust region and Kalman smoothing in sequential models (Särkkä and Svensson, 2020), we can use parallelized Kalman smoothers (Särkkä and García-Fernández, 2021; Chang et al., 2023) to achieve a parallel runtime that is logarithmic in the sequence length. We refer to this stabilized and parallelized Newton iteration for parallel RNN evaluation as “Evaluating Levenberg-Marquardt with Kalman smoothing”, or ELK. This name comes from damping Newton’s method with a trust region via the Levenberg-Marquardt algorithm, which can be evaluated in parallel over the sequence length using Kalman smoothing. We then combine quasi-Newton approximations with ELK, yielding an algorithm that is memory efficient, fast, and numerically stable, which we refer to as quasi-ELK.
We show empirically that quasi-DEER remains accurate, while also benefiting from reduced runtime and memory consumption. In regimes where DEER suffers from numerical instabilities and slow convergence, we show ELK and quasi-ELK enjoy fast, numerically stable convergence. We conclude by discussing the relative strengths and weaknesses of each method, and providing avenues for future research.
@@
Figure 1: Overview of the parallelizable methods we consider in this paper. Building on DEER (Lim et al., 2024), we introduce diagonal approximations to improve complexity (quasi-DEER, sec 4.1) and link to Kalman filtering and trust regions to improve stability (ELK, sec. 4.2). We combine these ideas in quasi-ELK (sec. 4.2).
@@
Table 1: Description of the relative strengths and weaknesses of the five evaluation methods we consider in this paper. We include discussion of this in Section 7.
@@
DEER (Lim et al., 2024) Yes $\mathcal{O}(TD^{3})$ $\mathcal{O}(TD^{2})$ Low
@@ 2 Problem Statement
[…]
@@
[…]
This residual can be interpreted as the one-step error incurred by assuming the $t^{\text{th}}$ state is $\mathbf{s}_{t}$ instead of $f(\mathbf{s}_{t-1})$. The solution of the state space model, $\mathbf{s}^{*}_{1:T}$, is the only trace with zero residual. Equivalently, it is the unique solution to the fixed point equation
@@
The conventional way of obtaining $\mathbf{s}^{*}_{1:T}$ is to apply $f$ sequentially $T$ times. Sequential evaluation always yields a valid trace, but it requires $\mathcal{O}(T)$ sequential operations (i.e., computational depth or span), and hence does not fully leverage the capabilities of parallel hardware. We aim to compute $\mathbf{s}^{*}_{1:T}$ in sublinear time using parallel computation.
@@
[…]
@@ 3 DEER: Newton’s Method for Parallel Evaluation of Sequential Models
Lim et al. (2024) propose DEER, an algorithm using Newton’s method for parallel evaluation of nonlinear sequential models, including both discrete-time nonlinear RNNs (GRUs, LSTMs, etc.) and neural ODEs (Chen et al., 2018; Kidger, 2021). In this paper, we focus on the discrete-time setting, and address natural questions that arise from Lim et al. (2024): how to scale Newton’s method, and how to make it numerically stable.
In this section we introduce DEER. We begin with a simplified derivation that emphasizes the link between Newton’s method on vector spaces and parallelizable linear recurrences. We then present a new proof that DEER theoretically always converges globally; but our proof also highlights why global convergence can be numerically unstable and/or slow in practice. We conclude by using these insights to discuss the weaknesses of DEER, and to motivate the methods we develop in Section 4.
@@ 3.1 Derivation of DEER from Newton’s Method
The original derivation of DEER used Newton’s method on Banach spaces and the Frécht derivative for continuous-time systems to derive the update and show convergence (Lim et al., 2024). We specialize to the setting of discrete-time RNNs, and present a streamlined derivation that more directly connects the structure of the Jacobian in eq. 3 to the linear recurrence relation in eq. 6. This connection highlights why DEER incurs cubic work in $D$, and may encounter numerical instabilities. We will also use this form to prove DEER’s global convergence in Section 3.2.
The $i^{\text{th}}$ Newton iterate for eq. 2, starting at $\mathbf{s}^{(i)}$, is given by
@@
Provided all of $\left\{\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right\}_{t=2}^{T}$ are finite, then the Jacobian defined in eq. 3 is invertible and all of the eigenvalues are equal to one. Storing and naively inverting the Jacobian is infeasible for large $D$ or $T$. However, since $J(\mathbf{s})$ is block bidiagonal, we can solve for $\Delta\mathbf{s}$ in eq. 5 by forward substitution. This reduces to a simple recursion with the initial condition $\Delta\mathbf{s}_{1}^{(i+1)}=-\mathbf{r}_{1}(\mathbf{s}^{(i)})$, and for $t>1$,
@@
DEER capitalizes on the fact that this recursion is linear, so it can be solved by a parallel associative scan (Lim et al., 2024; Blelloch, 1990; Smith et al., 2023). With $\mathcal{O}(T)$ processors, each Newton iteration can be performed in $\mathcal{O}(\log T)$ time.
@@ 3.2 Global Convergence of DEER
An open question is whether DEER converges globally for discrete-time RNNs. We prove that DEER does converge globally for discrete-time RNNs to the solution $\mathbf{s}^{*}_{1:T}$ of eq. 2 in at most $T$ steps.
@@
If the Jacobians of the dynamics function $f$ are finite everywhere in the state space, then undamped Newton’s method will converge to the true solution, $\mathbf{s}^{*}_{1:T}$, of the fixed point eq. 2 in at most $T$ Newton iterations, for any initial $\mathbf{s}^{(0)}_{1:T}$.
@@
The structure of $J(\mathbf{s})$ determines the recurrence in eq. 6. The update applied at time $t$, $\Delta\mathbf{s}_{t}^{(i+1)}$, from eq. 6 is the summation of a linearized $f$ applied to the update at time $t-1$, and the residual one-step error at time $t$. Therefore, if the previous timestep is correct (i.e. $\Delta\mathbf{s}_{t-1}^{(i+1)}=\mathbf{0}$), then the update at time $t$ is just the one-step residual, which is defined exactly as the error. Therefore, if the previous value is correct, the updated current value will be correct. Given that $f$ and $\mathbf{s}_{0}$ are fixed and known, the result follows that all $T$ timesteps will have zero residual after $T$ iterations. ∎
It is not immediately obvious from eq. 4 or the proof given by Lim et al. (2024) that DEER converges globally, but Proposition 1 shows that it in fact does, at least theoretically. This result has two crucial corollaries. First, after $i$ Newton iterations, $\mathbf{s}^{(i)}_{1:T}$ will have zero error for all $t\leq i$. Therefore, if the iteration encounters numerical instabilities (as Newton is prone to), we can simply use a heuristic of resetting $s_{t}^{(i)}$ to a finite value for all $t>i$. This heuristic preserves the solution for time indices $t\leq i$ and allows the optimization to continue, but it is equivalent to running Newton’s method from scratch on $\mathbf{s}_{i:T}$. This process is repeated until the entire trace has zero residual. A second corollary is that any set of finite matrices can replace $\left\{\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right\}_{t=2}^{T}$ in eq. 3 or eq. 6, and the resulting quasi-Newton method will still converge globally in at most $T$ iterations. This preservation of global convergence provides further motivation for exploring quasi-Newton methods, as we discuss in the next section.
@@ 3.3 Weaknesses of DEER
Despite the theoretical convergence of DEER, its formulation as a linear recurrence relation in eq. 6 highlights its limited scalability and stability. Scalability is limited because, in general, $\nicefrac{{\partial f}}{{\partial\mathbf{s}}}$ is a dense $D\times D$ matrix. Therefore the parallel associative scan, which uses matrix-matrix multiplications, has $\mathcal{O}(TD^{3})$ computational work and $\mathcal{O}(TD^{2})$ memory complexity. Stability is limited because we often have no control over the eigenvalues of $\nicefrac{{\partial f}}{{\partial\mathbf{s}}}$. If sufficiently many of these eigenvalues over the sequence length are larger in magnitude than one, then the linear recurrence relation will be numerically unstable. The heuristic approach of resetting unstable values is sufficient to ensure global convergence, but as we show in Section 6.3, it comes at the cost of runtime, as convergence is dramatically slower. These weaknesses motivate the development of two new techniques for parallelized evaluation of RNNs, quasi-DEER and ELK, which we discuss in the next section.

### v3

@@ Abstract
Transformers and linear state space models can be evaluated in parallel on modern hardware, but evaluating nonlinear RNNs appears to be an inherently sequential problem. Recently, however, Lim et al. [1] developed an approach called DEER, which evaluates nonlinear RNNs in parallel by posing the states as the solution to a fixed-point problem. They derived a parallel form of Newton’s method to solve the fixed-point problem and achieved significant speedups over sequential evaluation. However, the computational complexity of DEER is cubic in the state size, and the algorithm can suffer from numerical instability. We address these limitations with two novel contributions. To reduce the computational complexity, we apply quasi-Newton approximations and show they converge comparably to Newton, use less memory, and are faster. To stabilize DEER, we leverage a connection between the Levenberg-Marquardt algorithm and Kalman smoothing, which we call ELK. This connection allows us to stabilize Newton’s method while using efficient parallelized Kalman smoothing algorithms to retain performance. Through several experiments, we show that these innovations allow for parallel evaluation of nonlinear RNNs at larger scales and with greater stability.
@@ 1 Introduction
Parallel computation has helped fuel the rise of deep learning [2]. Architectures such as transformers [3] and linear RNNs [4, 5, 6, 7, 8] are specifically designed to allow parallelization over the length of the input sequence. However, most conventional nonlinear RNNs (e.g. Elman RNNs, GRUs [9], LSTMs [10] etc.) are not readily parallelizable over the sequence length due to their sequential architecture. Thus, they do not benefit as much from parallel hardware. Nonetheless, these nonlinear RNN architectures are still used widely across the scientific community [11, 12, 13]. Furthermore, recent work has suggested that linear RNNs (and transformers) are fundamentally limited in their expressivity compared to nonlinear RNNs [14]. Finally, nonlinear RNNs continue to be of significant interest in computational and theoretical neuroscience as models of neural systems [15, 16, 17, 18, 19, 20, 21, 22]. Therefore, scalable and stable parallelization methods for nonlinear RNNs offer significant benefits across many fields.
@@
Towards this goal, Lim et al. [1] proposed DEER, a method for evaluating a nonlinear RNN in parallel. DEER casts inference as finding the solution of a fixed-point equation designed specifically to capture the nonlinear dynamics of the RNN. Newton’s method is used to solve the resulting fixed-point equation. With good initialization, Newton’s method enjoys quadratic convergence rates [23, Chapter 11]. Lim et al. [1] also show that the inversion of the structured Jacobian matrix required by Newton’s method can be cast as an associative parallel scan [24]. DEER therefore reduces the evaluation runtime over sequential evaluation by as much as factor of twenty.
@@
However, DEER also inherits the weaknesses of Newton’s method and parallel scans. The first weakness is scalability. Let $D$ denote the state dimension and $T$ denote sequence length. Using a parallel scan to evaluate updates from Newton’s method, DEER inherits $\mathcal{O}(TD^{2})$ memory complexity and $\mathcal{O}(TD^{3})$ computational work [24]. These costs can be prohibitive in practical deep learning settings. The second limitation of DEER is numerical stability, inherited from Newton’s method. In general, undamped Newton’s method does not provide global convergence guarantees and in practice often diverges [23]. We seek to ameliorate both of these weaknesses.
To do this, we leverage two techniques: quasi-Newton approximations and trust regions. Quasi-Newton approximations are a common adaptation of Newton’s method, where approximate, but faster and less memory intensive updates are used instead of exact Newton steps. Empirically, these methods often expedite convergence in terms of wall-clock time, even though more iterations are performed. We apply quasi-Newton approximations to reduce the memory and compute required by DEER, and we find accelerated convergence and reduced memory consumption. Secondly, we leverage a connection between Newton’s method with a trust region and Kalman smoothing in sequential models [25]. This connection allows us to stabilize the Newton iteration by limiting the step size to the radius of the trust region, preventing large and numerically unstable steps. The update can be computed with a parallel Kalman smoother [26, 27], achieving a runtime that is logarithmic in the sequence length. We refer to DEER accelerated with a quasi-Newton approximation as quasi-DEER, and DEER stabilized with trust regions as “Evaluating Levenberg-Marquardt via Kalman” (ELK). We then combine these approaches to yield a fast and stable algorithm, which we call quasi-ELK.
Crucially, DEER, ELK, and their “quasi” variants are algorithms for parallelizing any discrete-time nonlinear dynamical system, including stateful architectures such as RNNs, that may or may not include stochasticity. We use “parallel” to refer to the fact that each iteration of our iterative algorithm operates on the entire $T$-length sequence (and not on each sequence element one at a time).
@@
Figure 1: Overview of the parallelizable methods we consider in this paper. We introduce diagonal approximations to improve complexity (quasi-DEER, Section 4.1) and link to Kalman filtering and trust regions to improve stability (ELK, Section 4.2). We combine these ideas in quasi-ELK (Section 4.2).
@@
Table 1: Description of the relative strengths and weaknesses of the five evaluation methods we consider. We include a discussion of this in Section 7.
@@
DEER [1] Yes $\mathcal{O}(TD^{3})$ $\mathcal{O}(TD^{2})$ Low
@@
We outline the key contributions and organization of the paper here. First, we introduce background material, particularly focusing on DEER [1], in Sections 2 and 3. We then present three short novel proofs: that DEER is globally convergent; that this convergence is robust to modifications of the linearized dynamics (Proposition 1); and that there is a unique solution with no local minima (Appendices A.1 and A.2). We then introduce quasi-Newton approximations to DEER to improve efficiency (quasi-DEER, Section 4.1), and trust regions to stabilize DEER (ELK, Section 4.2) We also provide an interpretation of how trust regions stabilize the dynamics by damping the eigenvalues of the Jacobians (Appendix A.3). We show empirically that quasi-DEER remains accurate, with reduced runtime and memory consumption (Section 6). In regimes where DEER is numerically unstable or convergences slowly, we show ELK and quasi-ELK can enjoy fast, numerically stable convergence. We conclude by discussing the relative strengths and weaknesses of each method, providing guidance on how to select and tune them, and highlighting avenues for future research (Section 7). We provide our code at https://github.com/lindermanlab/elk.
@@ 2 Problem Statement
We consider nonlinear Markovian state space models, with the state at time $t$ denoted $\mathbf{s}_{t}\in\mathbb{R}^{D}$ and nonlinear transition dynamics $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. We denote the full sequence of $T$ states as $\mathbf{s}_{1:T}\in\mathbb{R}^{T\times D}$. Note that we primarily focus on the transition dynamics, so we suppress any (possibly random) input dependence of the model in the notation. However, the algorithms in this paper extend easily to these situations.
@@
$\mathbf{r}(\mathbf{s}_{1:T}):=[\mathbf{s}_{1}-f(\mathbf{s}_{0}),\;\mathbf{s}_{2}-f(\mathbf{s}_{1}),\;\mathbf{s}_{3}-f(\mathbf{s}_{2}),\ldots,\mathbf{s}_{T}-f(\mathbf{s}_{T-1})]\in\mathbb{R}^{T\times D}.$ (1)
This residual can be interpreted as the one-step error incurred by assuming the $t^{\text{th}}$ state is $\mathbf{s}_{t}$ instead of $f(\mathbf{s}_{t-1})$. The solution of the state space model, $\mathbf{s}^{*}_{1:T}$, is the only trace with zero residual. Equivalently, it is the unique solution to the fixed-point equation
@@
The conventional way to obtain $\mathbf{s}^{*}_{1:T}$ is to apply $f$ sequentially $T$ times. Sequential evaluation always yields a valid trace, but it requires $\mathcal{O}(T)$ sequential operations (i.e. computational depth or span), and hence does not fully leverage the capabilities of parallel hardware. We aim to compute $\mathbf{s}^{*}_{1:T}$ in sublinear time using parallel computation.
@@
For notational brevity, we overload $\mathbf{s}$ and $\mathbf{r}$ to also denote vectors in $\mathbb{R}^{TD}$, representing flattened versions of $\mathbf{s}_{1:T}$ and $\mathbf{r}_{1:T}$. We can therefore write the Jacobian of the residual for the whole sequence, $J(\mathbf{s})$, as a $TD\times TD$ matrix with block bidiagonal structure of the form
@@ 3 DEER: Newton’s Method for Parallel Evaluation of Sequential Models
Lim et al. [1] propose DEER, an algorithm using Newton’s method for parallel evaluation of nonlinear sequential models, including both discrete-time nonlinear RNNs (GRUs, LSTMs, etc.) and neural ODEs [28, 29]. In this paper, we focus on the discrete-time setting, and address questions that arise from Lim et al. [1]: how to scale Newton’s method, and how to make it numerically stable.
In this section we introduce DEER. We begin with a simplified derivation that emphasizes the link between Newton’s method on vector spaces and parallelizable linear recurrences. We then present a new proof that DEER theoretically always converges globally. This proof also highlights why global convergence can be numerically unstable and/or slow in practice. We conclude by using these insights to discuss the weaknesses of DEER, and to motivate the methods we develop in Section 4.
@@ 3.1 Derivation of DEER from Newton’s Method
The original derivation of DEER used Newton’s method on Banach spaces and the Fréchet derivative for continuous-time systems to derive the update [1]. We specialize to the setting of discrete-time RNNs, and present a streamlined derivation that more directly connects the structure of the Jacobian in (3) to the linear recurrence relation in (6). This connection highlights why DEER incurs cubic work in $D$ and may encounter numerical instabilities. We will also use this form to prove DEER’s global convergence in Section 3.2.
The $i^{\text{th}}$ Newton iterate for (2), starting at $\mathbf{s}^{(i)}$, is given by
@@
Note this uses the root-finding view of Newton’s method, see Appendix C.2.
The Jacobian defined in (3) is invertible and all of the eigenvalues are equal to one. [^1] To see this fact, note that the characteristic polynomial of $J(\mathbf{s})$ in (3) is $(\lambda-1)^{TD}$. Storing and naively inverting the Jacobian is infeasible for large $D$ or $T$. However, since $J(\mathbf{s})$ is block bidiagonal, we can solve for $\Delta\mathbf{s}$ in (5) by forward substitution. This reduces to a simple recursion with the initial condition $\Delta\mathbf{s}_{1}^{(i+1)}=-\mathbf{r}_{1}(\mathbf{s}^{(i)})$, and for $t>1$,
@@
DEER uses the linearity of this recursion, solving it in parallel with a parallel associative scan [1, 5, 24]. Therefore, with $\mathcal{O}(T)$ processors, each Newton iteration can be performed in $\mathcal{O}(\log T)$ time.
We emphasize that the computation of the Newton step $\Delta\mathbf{s}$ in (5) is being parallelized. $J$ would, in general, be a $TD\times TD$ matrix that is prohibitive to store or invert. But by formulating this solve as an LDS in (6), we are able to parallelize the computation of $\Delta s$ (which consists of $T$ state updates, each of dimension $D$) over the sequence length. With sufficient processors, each update in (5) can be computed in $\mathcal{O}(\log T)$ time. Our implementation uses the parallel associative scan from JAX [30] (see Appendix B.6).
@@ 3.2 Global Convergence of DEER
We present a proof that DEER converges globally for discrete-time RNNs to the solution $\mathbf{s}^{*}_{1:T}$ of (2) in at most $T$ steps.
@@
Undamped Newton’s method will converge to the true solution, $\mathbf{s}^{*}_{1:T}$, of the fixed-point equation (2) in at most $T$ Newton iterations, for any initial $\mathbf{s}^{(0)}_{1:T}$.
@@
For the full proof by induction, see Appendix A.1. The structure of $J(\mathbf{s})$ determines the recurrence in (6). The update applied at time $t$, $\Delta\mathbf{s}_{t}^{(i+1)}$, from (6) is the summation of a linearized $f$ applied to the update at time $t-1$, and the residual one-step error at time $t$. Therefore, if the previous timestep is correct (i.e. $\Delta\mathbf{s}_{t-1}^{(i+1)}=\mathbf{0}$), then the update at time $t$ is just the one-step residual, which is defined exactly as the error. Therefore, if the previous value is correct, the updated current value will be correct. Given that $f$ and $\mathbf{s}_{0}$ are fixed and known, the result follows that all $T$ timesteps will have zero residual after $T$ iterations. ∎
It is not immediately obvious from (4) or past work Lim et al. [1] that DEER converges globally, but Proposition 1 shows that it does, at least theoretically. This result has two crucial corollaries. First, after $i$ Newton iterations, $\mathbf{s}^{(i)}_{1:T}$ will have zero error for all $t\leq i$. Therefore, if the iteration encounters numerical instabilities, as Newton is prone to, we can simply use a heuristic of resetting $s_{t}^{(i)}$ to a finite value for all $t>i$. This preserves the solution for time indices $t\leq i$ and allows the optimization to continue, but it is equivalent to running Newton’s method from scratch on $\mathbf{s}_{i:T}$. This process is repeated until the entire trace has zero residual. A second corollary is that any set of finite matrices can replace $\left\{\nicefrac{{\partial f}}{{\partial\mathbf{s}}}\right\}_{t=2}^{T}$ in (3) or (6), and the resulting quasi-Newton method will still converge globally in at most $T$ iterations. This preservation of global convergence provides further motivation for exploring quasi-Newton methods, as we discuss in the next section.
@@ 3.3 Weaknesses of DEER
Despite the theoretical convergence of DEER, its formulation as a linear recurrence relation in (6) highlights its limited scalability and stability. Scalability is limited because, in general, $\nicefrac{{\partial f}}{{\partial\mathbf{s}}}$ is a dense $D\times D$ matrix. Therefore the parallel associative scan, which uses matrix-matrix multiplications, has $\mathcal{O}(TD^{3})$ computational work and $\mathcal{O}(TD^{2})$ memory complexity. Stability is limited because we often have no control over the eigenvalues of $\nicefrac{{\partial f}}{{\partial\mathbf{s}}}$. If sufficiently many of these eigenvalues over the sequence length are larger in magnitude than one, then the linear recurrence relation will be numerically unstable. The heuristic approach of resetting unstable values is sufficient to ensure global convergence, but as we show in Section 6.3, it comes at the cost of runtime, as convergence is dramatically slower. These weaknesses motivate the development of two new techniques for parallel evaluation of RNNs: quasi-DEER and ELK, which we discuss in the next section.